OpenStudy (gandalf123):

The force of gravity, F, exerted between two objects is equal to the product of the gravitational constant, G, the mass of the first object, m1, and the mass of the second object, m2, divided by the square of the distance between their centers, d. This is often used to determine the gravitational attraction between two massive bodies, such as planets, in space.

OpenStudy (gandalf123):

b. If you were a physicist trying to determine the gravitational constant and you were able to measure the force of gravity exerted between two objects, their masses, and the distance between their centers, how would you rewrite this formula?

OpenStudy (jadzia):

Universal Gravitational Law is: \(\sf \Large F = G \frac{m_im_2}{d^2}\) Evaluate the formula in terms of G: 1. Multiply both sides by \(\sf d^2\) 2. Divide both sides by \(\sf m_1m_2\) 3. Simplify and you'll get it.

OpenStudy (rushwr):

So the end result would be \[\frac{ F * d ^{2} }{ m _{1}m _{2} } = G\]